OpenStudy (anonymous):

Solve: 2 6 - = - 7 x

OpenStudy (anonymous):

A. x = 21 B. x = 11 C. x =12/7 D. x =2/7

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

@brookie5150

OpenStudy (anonymous):

2/7 = 6/x 2x =42 x =21

OpenStudy (anonymous):

c:

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yup